# Synthesizing Programmatic Reinforcement Learning Policies with Large Language Model Guided Search

### v1

@@ Appendix C Karel tasks
There are five knowledge we tell the Large Language Model: task name, task map, task agent position, task goal, and task return. Here is all the information about these ten tasks. Appendix F show how these five different components are filled into the placeholders in the user prompt. There are two sets of tasks: Karel [57] and Karel-Hard [37]. Section C.1 lists all six Karel tasks, and Section C.2 lists all four Karel-Hard tasks. There are six Karel tasks: StairClimber, FourCorners, TopOff, Maze, Harvester, and CleanHouse. Example figures of Karel tasks are in Figure 9. On the other hand, four of the Karel-Hard tasks are DoorKey, OneStroke, Seeder, and Snake. Figure 10(d) shows the initial and desired states of the Karel-Hard tasks.
@@ C.1.1 The task prompt of StairClimber
Knowledge Prompt Text
@@ C.1.2 The task prompt of Maze
Knowledge Prompt Text
@@ C.1.3 The task prompt of FourCorners
Knowledge Prompt Text
@@ C.1.4 The task prompt of TopOff
Knowledge Prompt Text
@@ C.1.5 The task prompt of Harvester
Knowledge Prompt Text
@@ C.1.6 The task prompt of CleanHouse
Knowledge Prompt Text
@@ C.2 Karel-Hard tasks
Figure 10: Illustrations of the initial and final state of each task in the Karel-Hard problem set introduced in by Liu et al. [37]. The position of markers, walls, and agent’s position are randomly set according to the configurations of each tasks.
@@ C.2.1 The task prompt of DoorKey
Knowledge Prompt Text
@@ C.2.2 The task prompt of OneStroke
Knowledge Prompt Text
@@ C.2.3 The task prompt of Seeder
Knowledge Prompt Text
@@ C.2.4 The task prompt of Snake
Knowledge Prompt Text
@@
Appendix D Baselines and their hyperparameters
We compares to four baselines proposed by Trivedi et al. [57],  Liu et al. [37], and Carvalho et al. [10]. LEAPS, HPRL, and CEBS search programs in the latent spaces. HC search program in the programmatic space. The pseudo-code of CEBS is described in Algorithm 1, and the pseudo-code HC is described in Algorithm 2.
D.1 Learning Embeddings for Latent Program Synthesis (LEAPS)
The hyperparameters of the LEAPS baseline are from LEAPS [57] and HPRL [37]. We downloaded the pre-trained weights of LEAPS and used the parameters for searching programs. We do not use the official implementation of LEAPS, thus we do not have hyperparameters of exponential decay. Also, the unofficial implementation samples candidate from a fixed normal distribution which is hyperparameter in LEAPS [57]. In  Table LABEL:table:leaps_karel_setting, $K$ is the neighborhood size and $E$ is the candidate, $\sigma$ is the noise scale. In LEAPS and HPRL, they use an elite ratio to represent the candidate programs. For better understanding, we round $E$ to be an integer.
Table 7: LEAPS hyperparameter settings on all Karel tasks.
Task Name $K$ $\sigma$ $E$
StairClimber 32 0.25 2
FourCorners 64 0.5 13
TopOff 64 0.25 3
Maze 16 0.1 2
CleanHouse 32 0.25 2
Harvester 32 0.5 3
DoorKey 32 0.25 3
OneStroke 64 0.5 3
Seeder 32 0.25 3
Snake 32 0.25 6
D.2 Hierarchical programmatic reinforcement learning (HPRL)
HPRL [37] aims to improve LEAPS by composing several programs to represent more complex behaviors. Given a learned latent space, meta-policy learns to predict a sequence of actions, i.e. programs, by optimizing the return obtained by executing the composed programs using reinforcement learning. HPRL considers the discount factor $\gamma$ in the meta-MDP; as a result, the evaluation method is not the same as our problem formulation. To ensure a fair comparison, we record one million programs explored by meta-policy in the training stage. In HRPL, the authors use the training step as a hyperparameter. One training step is one step of meta-policy, and the maximum episode length is set to five as the original setting in the HPRL paper. In Table LABEL:table:hprl_karel_setting, we list the hyperparameters we modified in the HPRL training script. All other hyperparameters remain the same as described in HPRL.
Table 8: HRPL hyperparameter settings on all Karel tasks.
Task Name training steps height weight
StairClimber 50K 12 12
FourCorners 500K 12 12
TopOff 5M 12 12
Maze 50K 8 8
CleanHouse 500K 12 22
Harvester 5M 8 8
DoorKey 5M 8 8
OneStroke 5M 8 8
Seeder 5M 8 8
Snake 5M 8 8
D.3 Cross-entropy beam search (CEBS)
CEBS [10] also utilizes the pre-trained VAE weight from LEAPS [57] to search program in the latent space. All the hyperparameter follows the original CEBS with neighborhood $K$ equal to 64, candidate $E$ equal to 16, and noise ratio $\sigma$ equal to 0.25 for all ten Karel tasks. The pseudo-code is described in  Algorithm 1.
D.4 Hill climbing (HC)
HC [10] is a state-of-the-art algorithm solving Programmatic Reinforcement Learning tasks. All the hyperparameter follows the original HC with neighborhood $K$ equal to 250 for all ten Karel tasks. The pseudo-code is described in  Algorithm 2.
Algorithm 1 Cross-entropy beam search algorithm  [10]
0:  $\rho^{*}$ the highest averaged return over 32 task variants.
1:  $z\sim N(0,\mathbf{I})$
2:  $\rho\leftarrow\text{VAE.decode}(z)$
3:  $steps\leftarrow 0$
4:  $Return\leftarrow\text{evaluate}(\rho,T)$
5:  $\rho^{*}\leftarrow\rho$
6:  $Mean\leftarrow-\infty$
7:  $P\leftarrow\text{get-neighbor}(\rho,K)$
8:  while $steps<1000000$ do
9:     $Candidates\leftarrow\text{[]}$]
10:     for each $\rho_{\text{new}}$ in $P$ do
11:        $r\leftarrow\text{evaluate}(\rho_{\text{new}},T)$
12:        $Candidates.\text{append}(r,\rho_{\text{new}})$
13:        if $r>Return$ then
14:           $\rho^{*}\leftarrow\rho_{\text{new}}$
15:           $Return\leftarrow r$
16:        end if
17:     end for
18:     $Elites\leftarrow\text{Top-E}(Candidates)$
19:     if $Mean>\text{get-mean}(Elites)$ then
20:        break
21:     end if
22:     if $Return=1$ then
23:        break
24:     end if
25:     $Mean\leftarrow\text{get-mean}(Elites)$
26:     $P\leftarrow\text{[]}$]
27:     for  each $\rho_{\text{new}}$ in $Elites$ do
28:        $P.\text{extend}(\text{get-neighbor}(\rho_{\text{new}},K/E))$
29:     end for
30:     $steps\leftarrow steps+K$
31:  end while
Algorithm 2 Hill climbing algorithm  [10]
0:  $T$, the task; $K$, number of neighborhood.
0:  $\rho^{*}$, the highest averaged return over 32 task variants.
1:  Initialize $\rho$ with a random solution
2:  $Return\leftarrow\text{evaluate}(\rho,T)$
3:  $\rho^{*}\leftarrow\rho$
4:  $improved\leftarrow\text{True}$
5:  $steps\leftarrow 0$
6:  while $improved$ $steps<n$ do
7:     $improved\leftarrow\text{False}$
8:     $Neighbors\leftarrow\text{get-neighbor}(\rho,$K))))
9:     for each $\rho_{\text{new}}$ in $Neighbors$ do
10:        $r_{\text{new}}\leftarrow\text{evaluate}(\rho_{\text{new}},T)$
11:        if $r_{\text{new}}>Return$ then
12:           $\rho^{*}\leftarrow\rho_{\text{new}}$
13:           $Return\leftarrow r_{\text{new}}$
14:           $improved\leftarrow\text{True}$
15:           break
16:        end if
17:        $steps\leftarrow steps+1$
18:     end for
19:     $\rho\leftarrow\rho^{*}$
20:  end while
Appendix E An example pipeline of our method
This is our pipeline of getting programs from LLM with the task DoorKey.  Section E.1 details the system prompt,  Section E.2 provides all the information in one specific task,  Section E.3 lists the LLM response, and  Section E.4 gives an example how we get the Karel DSL program.
E.1 The system prompt
@@
E.2 The user prompt of the task DoorKey
@@
E.3 The sample response of LLM in the task DoorKey
@@
E.4 Apply post-processing and retrieve the final program
For the Python program part in LLM response, we convert the Python program to a DSL program through the rule-based method and apply post-processing to correct minor mistakes. For the DSL part, we directly apply post-processing to correct minor mistakes. Section E.4.1 presents the original Python program, the converted DSL, and the post-processed DSL. Section E.4.2 shows the original DSL and the post-processed DSL. All of the post-processing rules are listed in Table 5.
E.4.1 Python program part
@@ E.4.2 DSL program part
Appendix F LLM prompts for ablations and revision
@@
Ablation prompts for program generation methods. We conduct an ablation study in Section 5.3 to justify that our LLM-generating program method performs best in acceptance rate and best return. Here are the complete prompts. There are three approaches to generating DSL programs: Pythonic-DSL, Python, and DSL. There are two types of prompt, system prompt and user prompt, in all approaches. In the system prompt, both the Python and the Pythonic-DSL approaches contain limitations of Python usage. The DSL approach contains the grammar of the Karel, and the Pythonic-DSL contains the paired Python-like and Karel production rules. All of the system prompts contain the environment physics, action, and perception. On the other hand, the user prompt contains five placeholders for task name, map position, initial position, task goal, and task reward. In Appendix C, all of the task-dependent information can be filled in the placeholders. Section E.2 shows the full Pythonic-DSL prompt with the task DoorKey filled in the placeholders. The Pythonic-DSL prompts are in  Section F.1, and the Python prompts are in Section F.2, and the DSL prompts are in Section F.3.
@@
Prompts for LLM revision. We list all of the user prompts in the experiment of LLM revision, the system prompt is the same as the one in Section F.1. There are four approaches to revising the program: Regenerate, Regenerate with reward, Agent execution trace, and Agent and program execution trace. We ask the LLM to regenerate with all generated programs in the last round without repetition, and the Regenerate prompt is in Section F.4. We ask the LLM to regenerate with the program relating to reward, and the revision prompt of the Regenerate with reward is in Section F.5. We utilize the execution trace of the Karel agent in the grid world, and the revision prompt of the Agent execution trace is in Section F.6. We provided the action/perception call and executing line, and the revision prompt of the Agent and program execution trace is in Section F.7. The program Figure 11 is the revision target of our method revision method in Section F.6 and Section F.7. Figure 11 is the trajectories of the program we used in method Section F.6 and Section F.7. In the example, the origin program has an average return of 0.5. LLM revision with agent trace can reach 0.640625, and LLM revision with both agent and program traces can reach a great result of 0.8125.
@@
[!ht] {minted}python def run(): while noMarkersPresent(): if frontIsClear(): move() else: turnLeft() pickMarker() while frontIsClear(): move() turnRight() while noMarkersPresent(): if frontIsClear(): move() else: turnLeft() if frontIsClear(): move() turnRight() putMarker() The revision results of Section F.6.
@@
[!ht] {minted}python def run(): while noMarkersPresent(): if frontIsClear(): move() else: turnLeft() pickMarker() while noMarkersPresent(): if rightIsClear(): turnRight() move() else: if frontIsClear(): move() else: turnLeft() putMarker() The revision results of Section F.7.
F.1 Pythonic-DSL
F.1.1 System prompt

### v2

@@ Appendix C Karel tasks
We give the large language model 5 pieces of information related to the tasks: task name, map description, initial Position, task goal, and task return. By converting the original task description into these 5 categories, this design allows the user to easily fill in all categories for new tasks. Here we list all the information pieces for these tasks. Section I.1.2 shows how these 5 different components are filled into the placeholders in the user prompt. Section D.2 show all filled results with the task DoorKey. There are three sets of tasks: Karel [Trivedi et al., 2021], Karel-Hard [Liu et al., 2023], and new Karel tasks. Section C.1 lists all 6 Karel tasks, Section C.2 lists all 4 Karel-Hard tasks, and Section C.3 lists the 2 new tasks. Example figures of Karel tasks are in Figure 10, Karel-Hard tasks are in Figure 11 and new Karel tasks are in Figure 12.
@@ C.1.1 The task prompt of StairClimber
Purpose Prompt Text
@@ C.1.2 The task prompt of Maze
Purpose Prompt Text
@@ C.1.3 The task prompt of FourCorners
Purpose Prompt Text
@@ C.1.4 The task prompt of TopOff
Purpose Prompt Text
@@ C.1.5 The task prompt of Harvester
Purpose Prompt Text
@@ C.1.6 The task prompt of CleanHouse
Purpose Prompt Text
@@ C.2 Karel-Hard tasks
Figure 11: Illustrations of the initial and final state of each task in the Karel-Hard problem set introduced in by Liu et al. [2023]. The position of markers, walls, and agent’s position are randomly set according to the configurations of each task.
@@ C.2.1 The task prompt of DoorKey
Purpose Prompt Text
@@ C.2.2 The task prompt of OneStroke
Purpose Prompt Text
@@ C.2.3 The task prompt of Seeder
Purpose Prompt Text
@@ C.2.4 The task prompt of Snake
Purpose Prompt Text
@@
C.3 New Karel tasks
To showcase the extensibility of our proposed LLM-GS framework, we additionally propose two novel new tasks, PathFollow and WallAvoider. Here we list the task prompts.
(a) PathFollow
(b) WallAvoider
Figure 12: Illustrations of the initial and final state of the two new proposed tasks.
C.3.1 The task prompt of PathFollow
Purpose Prompt Text
Task Name PATHFOLLOW
Map Description The map is given by a 8x8 grid surrounded by walls. There is a rugged ascending markers line that starts from the bottom left cell and randomly grows either north or to the east until it reaches the top right cell. Resulting in a rugged markers line connecting the bottom left cell and the top right cell.
Initial Position The agent starts on the bottom left cell of the map facing north.
Task Goal The goal of the agent is to collect every marker on that rugged markers line without leaving the rugged markers line two cells away.
Task Reward Return is given by the number of picked-up markers divided by the total number of markers. Placing any marker or leaving the rugged markers line two cells away will have a negative return as -1.0 and terminate the episode.
C.3.2 The task prompt of WallAvoider
Purpose Prompt Text
Task Name WALLAVOIDER
Map Description The map is given by an empty 8x5 grid surrounded by walls.
Initial Position The agent starts on a random cell of the map facing random directions.
Task Goal The goal of the agent is to place exactly one marker in every interior cell of the map, which refers to the cells that are not adjacent to any wall.
Task Reward Return is given by the number of interior cells with exactly one marker divided by the total number of interior cells. Picking up the marker, putting more than one marker on one cell, or putting any marker on the cell adjacent to any wall will terminate the episode with a -1 return.
Appendix D An example pipeline of our method
This is our pipeline of getting programs from LLM with the task DoorKey. Section D.1 details the system prompt, Section D.2 provides all the information in the task DoorKey. Section D.3 lists the LLM response, and Section D.4 gives an example of how we get the Karel DSL program.
D.1 The system prompt
@@
D.2 The user prompt of the task DoorKey
@@
D.3 The sample response of LLM in the task DoorKey
@@
D.4 Apply post-processing and retrieve the LLM-initialized programs
For the Python program part in LLM response, we convert the Python program to a DSL program through the rule-based method and apply post-processing to correct minor mistakes. For the DSL part, we directly apply post-processing to correct minor mistakes. Section D.4.1 presents the original Python program, the converted DSL, and the post-processed DSL. Section D.4.2 shows the original DSL and the post-processed DSL. All of the post-processing rules are listed in Table 6.
D.4.1 Python program part
@@ D.4.2 DSL program part
Appendix E Scheduled Hill Climbing detail
This Scheduler is designed for improving sample efficiency. $n$ is the number of programs evaluated, and $K_{start}$ and $K_{end}$ are hyper-parameters representing the initial and terminal number of neighbors to search, respectively. $r(n)$ is a sinusoidal function that smoothly increase from 00 to $1$ over the course of $N$ total programs evaluated in the environment. We provide detailed equation in Equation 2 and Equation 3
$log_{2}{k(n)}$ $=(1-r(n))\log_{2}{K_{start}}+r(n)\log_{2}{K_{end}}\text{,}$ (2)
Appendix F Baselines and their hyperparameters
We compare our framework to four baselines, LEAPS [Trivedi et al., 2021], HPRL [Liu et al., 2023], CEBS [Carvalho et al., 2024], and HC [Carvalho et al., 2024]. LEAPS, HPRL, and CEBS search programs in the latent spaces. HC search program in the programmatic space. The pseudo-code of CEBS is described in Algorithm 1, and the pseudo-code of HC is described in Algorithm 2.
F.1 LEAPS
The hyperparameters of the LEAPS baseline are from LEAPS [Trivedi et al., 2021] and HPRL [Liu et al., 2023]. We downloaded the pre-trained weights of LEAPS and used the parameters for searching programs. We use the re-implementation of LEAPS from Carvalho et al. [2024], thus we do not have hyperparameters of exponential decay. Also, the re-implementation samples candidate from a fixed normal distribution which is hyperparameter in LEAPS [Trivedi et al., 2021]. In Table LABEL:table:leaps_karel_setting, $K$ is the neighborhood size and $E$ is the candidate, $\sigma$ is the noise scale. In LEAPS and HPRL, they use an elite ratio to represent the candidate programs. For better understanding, we round $E$ to be an integer.
Table 7: LEAPS hyperparameter settings on all Karel tasks.
Task Name $K$ $\sigma$ $E$
StairClimber 32 0.25 2
FourCorners 64 0.5 13
TopOff 64 0.25 3
Maze 16 0.1 2
CleanHouse 32 0.25 2
Harvester 32 0.5 3
DoorKey 32 0.25 3
OneStroke 64 0.5 3
Seeder 32 0.25 3
Snake 32 0.25 6
F.2 Hierarchical programmatic reinforcement learning (HPRL)
HPRL [Liu et al., 2023] aims to improve LEAPS by composing several programs to represent more complex behaviors. Given a learned latent space, meta-policy learns to predict a sequence of actions, i.e. programs, by optimizing the return obtained by executing the composed programs using reinforcement learning. HPRL considers the discount factor $\gamma$ in the meta-MDP; as a result, the evaluation method is not the same as our problem formulation. To ensure a fair comparison, we record one million programs explored by meta-policy in the training stage. In HRPL, the authors use the training step as a hyperparameter. One training step is one step of meta-policy, and the maximum episode length is set to 5 as the original setting in the HPRL paper. In Table LABEL:table:hprl_karel_setting, we list the hyperparameters we modified in the HPRL training script. All other hyperparameters remain the same as described in HPRL.
Table 8: HRPL hyperparameter settings on all Karel tasks.
Task Name training steps height weight
StairClimber 50K 12 12
FourCorners 500K 12 12
TopOff 5M 12 12
Maze 50K 8 8
CleanHouse 500K 12 22
Harvester 5M 8 8
DoorKey 5M 8 8
OneStroke 5M 8 8
Seeder 5M 8 8
Snake 5M 8 8
F.3 Cross-entropy beam search (CEBS)
CEBS [Carvalho et al., 2024] also utilizes the pre-trained VAE weight from LEAPS [Trivedi et al., 2021] to search program in the latent space. All the hyperparameter follows the original CEBS with neighborhood $K$ equal to 64, candidate $E$ equal to 16, and noise ratio $\sigma$ equal to 0.25 for all ten Karel tasks. The pseudo-code is described in Algorithm 1.
F.4 Hill climbing (HC)
HC [Carvalho et al., 2024] is a state-of-the-art algorithm solving Programmatic Reinforcement Learning tasks. All the hyperparameter follows the original HC with neighborhood $K$ equal to 250 for all ten Karel tasks. The pseudo-code is described in Algorithm 2.
Algorithm 1 Cross-entropy beam search algorithm  [Carvalho et al., 2024]
0:  $\rho^{*}$ the highest averaged return over 32 task variants.
1:  $z\sim N(0,\mathbf{I})$
2:  $\rho\leftarrow\text{VAE.decode}(z)$
3:  $steps\leftarrow 0$
4:  $Return\leftarrow\text{evaluate}(\rho,T)$
5:  $\rho^{*}\leftarrow\rho$
6:  $Mean\leftarrow-\infty$
7:  $P\leftarrow\text{get-neighbor}(\rho,K)$
8:  while $steps<1000000$ do
9:     $Candidates\leftarrow\text{[]}$]
10:     for each $\rho_{\text{new}}$ in $P$ do
11:        $r\leftarrow\text{evaluate}(\rho_{\text{new}},T)$
12:        $Candidates.\text{append}(r,\rho_{\text{new}})$
13:        if $r>Return$ then
14:           $\rho^{*}\leftarrow\rho_{\text{new}}$
15:           $Return\leftarrow r$
16:        end if
17:     end for
18:     $Elites\leftarrow\text{Top-E}(Candidates)$
19:     if $Mean>\text{get-mean}(Elites)$ then
20:        break
21:     end if
22:     if $Return=1$ then
23:        break
24:     end if
25:     $Mean\leftarrow\text{get-mean}(Elites)$
26:     $P\leftarrow\text{[]}$]
27:     for  each $\rho_{\text{new}}$ in $Elites$ do
28:        $P.\text{extend}(\text{get-neighbor}(\rho_{\text{new}},K/E))$
29:     end for
30:     $steps\leftarrow steps+K$
31:  end while
Algorithm 2 Hill climbing algorithm  [Carvalho et al., 2024]
0:  $T$, the task; $K$, number of neighborhood.
0:  $\rho^{*}$, the highest averaged return over 32 task variants.
1:  Initialize $\rho$ with a random solution
2:  $Return\leftarrow\text{evaluate}(\rho,T)$
3:  $\rho^{*}\leftarrow\rho$
4:  $improved\leftarrow\text{True}$
5:  $steps\leftarrow 0$
6:  while $improved$ and $steps<n$ do
7:     $improved\leftarrow\text{False}$
8:     $Neighbors\leftarrow\text{get-neighbor}(\rho,$K))))
9:     for each $\rho_{\text{new}}$ in $Neighbors$ do
10:        $r_{\text{new}}\leftarrow\text{evaluate}(\rho_{\text{new}},T)$
11:        if $r_{\text{new}}>Return$ then
12:           $\rho^{*}\leftarrow\rho_{\text{new}}$
13:           $Return\leftarrow r_{\text{new}}$
14:           $improved\leftarrow\text{True}$
15:           break
16:        end if
17:        $steps\leftarrow steps+1$
18:     end for
19:     $\rho\leftarrow\rho^{*}$
20:  end while
Appendix G Data Leakage
While our proposed framework shows significant improvement over the baselines, some may question if this improvement only comes from the LLM “memorizing” all the answers. Indeed, some ground truth solutions have been documented in the previous literature [Trivedi et al., 2021, Liu et al., 2023, Carvalho et al., 2024], thus making data leakage a potential concern. Still, we would like to rule out this possibility from three aspects: the timeline of previous works, the LLMs’ understanding of the Karel tasks, and the innovation of two novel tasks. Thus proving our framework utilizes the LLMs’ understanding and reasoning ability not just its internal knowledge.
Timeline. For the Karel-Hard problem set introduced by Liu et al. [2023] in July 2023, the optimal programs were not included in Liu et al. [2023] since their proposed method could only partially solve the tasks. Carvalho et al. [2024], made public in 2024, are the first to provide optimal programs for the Karel-Hard tasks. Hence, our LLM (gpt-4-turbo-2024-04-09 with knowledge up to December 2023) could not access the optimal Karel-Hard tasks programs.
Does the LLM know the Karel tasks?. To verify if the LLM knows the Karel tasks used in our work, we prompt GPT-4 by providing the paper names of LEAPS [Trivedi et al., 2021] and HPRL [Liu et al., 2023] and ask GPT-4 to list all the names and descriptions of the Karel tasks used in these works. Among all the responses generated by GPT-4, the only task closely resembling the Karel tasks used in our work is ”Maze Solving.” However, all other tasks described in the GPT-4’s response are drastically distinct from the ones we used in our work. The full prompt and response are listed in Figure 13 and Figure 14.
Two novel tasks. We additionally design two novel tasks that have not been used in any existing work, PathFollow, and WallAvoider, as illustrated in Figure 12. The details of these two novel tasks are in Section C.3. The results in Figure 8 show that our framework can extend to new tasks that are impossible to leak to LLMs.
{mdframed}
The Karel environment is used to test Programmatic Reinforcement Learning (PRL) algorithms and includes a variety of tasks. There are six tasks in the Karel test set and four more challenging tasks in the Karel-hard test set. These tasks are utilized in the research papers ”Learning to Synthesize Programs as Interpretable and Generalizable Policies” and ”Hierarchical Programmatic Reinforcement Learning via Learning to Compose Programs.”
Can you list all ten tasks and their task descriptions based on what you have learned?
Figure 13: The prompt that asks GPT-4 about the Karel PRL tasks
{mdframed}
Sure! The Karel environment, based on the educational programming language designed to introduce students to the basics of programming, provides programmatic challenges that both regular and hierarchical reinforcement learning algorithms can address. Here is a summary of the tasks typically included in both the Karel test set and the Karel-hard set based on common challenges noted in similar environments and research contexts:
Karel Test Set Tasks
  1. 1.
    Move to Wall: The task is for the agent (Karel) to move forward until it encounters a wall.
  2. 2.
    Turn Right: Teach Karel to turn right using a sequence of commands (since Karel inherently only has a command to turn left).
  3. 3.
    Collect Marker: Karel needs to navigate towards and collect a marker from the grid.
  4. 4.
    Avoid Hurdles: Karel must move from one point to another while avoiding hurdles placed in its path.
  5. 5.
    Put Marker and Move: This task combines placing a marker in a specified location and then moving to a different location.
  6. 6.
    Maze Solving: Karel needs to find a path from one point to another in a maze-like structure.
Karel-Hard Test Set Tasks
  1. 1.
    Multiple Markers: Karel might need to collect or place multiple markers in various locations, requiring more complex navigation and memory usage.
  2. 2.
    Dynamic Obstacles: The environments where obstacles may change location or appear suddenly, requiring Karel to adapt its path dynamically.
  3. 3.
    Conditional Logic Usage: Tasks where Karel must use sensors to make decisions based on the conditions of its environment (e.g., If front is clear, move forward; otherwise, turn left).
  4. 4.
    Complex Maze: More intricate and challenging mazes that require advanced pathfinding algorithms and possibly backtracking strategies.
While these are typical tasks deduced from the context of using Karel for testing PRL algorithms, please note that specific details or task names might vary based on the particular implementation or the focuses of different research papers or setups within the ”Karel” environment. Always refer back to the specific research documents or source materials for the most accurate descriptions and objectives of these tasks.
Figure 14: The GPT-4 response about Karel PRL tasks
Appendix H Aggregated performance across all tasks
To clearly examine the performance across all the tasks, we include a plot that aggregates the performance of all main results (Figure 4) in Figure 15. Our proposed LLM-GS framework significantly outperforms the existing methods.
Figure 15: Aggregated performance across all the tasks. The aggregate performance across all ten tasks. Our proposed LLM-GS framework surpasses all existing methods by a large margin.
Appendix I LLM prompts for ablations and revision
@@
Ablation prompts for program generation methods. We conduct an ablation study in Section 5.3 to justify that our LLM-generating program method performs best in acceptance rate and best return. Here we list the complete prompts. There are three approaches to generating DSL programs: Pythonic-DSL, Python, and DSL. There are two types of prompt, system prompt and user prompt, in all approaches. In the system prompt, both the Python and the Pythonic-DSL approaches contain limitations of Python usage. The DSL approach contains the grammar of the Karel, and the Pythonic-DSL contains the paired Python-like and Karel production rules. All of the system prompts contain the environment physics, action, and perception. On the other hand, the user prompt contains five placeholders for task name, map description, initial position, task goal, and task reward. In Appendix C, all of the task-dependent information can be filled in the placeholders. Section D.2 shows the full Pythonic-DSL prompt with the task DoorKey filled in the placeholders. The Pythonic-DSL prompts are in Section I.1, and the Python prompts are in Section I.2, and the DSL prompts are in Section I.3.
@@
Prompts for LLM revision. We list all of the user prompts in the experiment of LLM revision, the system prompt is the same as the one in Section I.1. We implement four approaches to revising the program: Regenerate, Regenerate with reward, Agent execution trace, and Agent and program execution trace. We ask the LLM to regenerate with all generated programs in the last round without repetition, and the Regenerate prompt is in Section I.4. We ask the LLM to regenerate with the program relating to reward, and the revision prompt of the Regenerate with reward is in Section I.5. We utilize the execution trace of the Karel agent in the grid world, and the revision prompt of the Agent execution trace is in Section I.6. We provided the action/perception call and executing line, and the revision prompt of the Agent and program execution trace is in Section I.7. The program Figure 16 is the revision target of our method revision method in Section I.6 and Section I.7. Figure 16 is the trajectories of the program we used in method Section I.6 and Section I.7. In the example, the origin program has an average return of 0.5. LLM revision with agent trace can reach 0.640625, while LLM revision with both agent and program traces can reach a result of 0.8125.
@@
[!ht] {minted}python def run(): while noMarkersPresent(): if frontIsClear(): move() else: turnLeft() pickMarker() while frontIsClear(): move() turnRight() while noMarkersPresent(): if frontIsClear(): move() else: turnLeft() if frontIsClear(): move() turnRight() putMarker() The revision results of Section I.6.
@@
[!ht] {minted}python def run(): while noMarkersPresent(): if frontIsClear(): move() else: turnLeft() pickMarker() while noMarkersPresent(): if rightIsClear(): turnRight() move() else: if frontIsClear(): move() else: turnLeft() putMarker() The revision results of Section I.7.
I.1 Pythonic-DSL
I.1.1 System prompt